C= 450 +0.4Y
I=350
G=150
X=70
Z= 35+0.1Y
T= 0.15Y
Yf= 1550

Calculate the tax revenue to the government of this country when this economy remains in equilibrium
1
Expert's answer
2020-11-30T16:35:19-0500

 Aggregate expenditure function, "AE=C+I +G +X-IM(importst)"


"C= 450 +0.4Yd"


Where Yd = Y- T


"AE= 450 +0.4(Y-T)+ 350 +150 +70- 0"


"Y=920 +0.4Y-0.06"


"Y-0.4Y=920-0.06"


"0.6Y= 919.94"


"0.6Y=919.94"

"Y=919.94\div0.6 = 1533.23"